Richetti Band Uniforms uses a standard costing system. The standard material and labor costs for producing a marching band hat are as follows:

Materials (0.90 yards × $11.70) $10.53
Direct labor (1.20 hours × $13.30) $15.96


During May, the company produced 3,770 band hats; 4,400 yards of material were purchased for $42,240, and 3,270 yards of material were used in production. Also during May, 4,790 direct labor hours were worked at a cost of $57,480. Calculate material price and quantity variances and labor rate and efficiency variances. Indicate whether the variances are favorable or unfavorable. (Round intermediate calculations to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 125. Enter all variances as a positive number.)

Material Price Variance $

FavorableNeither Unfavorable nor FavorableUnfavorable

Material Quantity Variance $

FavorableNeither Unfavorable nor FavorableUnfavorable

Labor Rate Variance $

Neither Unfavorable nor FavorableUnfavorableFavorable

Labor Efficiency Variance $

Neither Unfavorable nor FavorableUnfavorableFavorable
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (11.70 - 42240/4400) * 4400

= 9240 Favourable

Materials Quantity Variance = (SQ - AQ) * SP

= (3770*0.90 - 3270)*11.70

= 1439 Favourable

Labor Rate Variance = (SR-AR) *AH

= (13.30 - 57480/4790),4790

= 6227 Favourable

Labor efficiency variance = (SH-AH) * SR

= (3770*1.20 - 4790)*13.30

= 3538 Unfavorable

Note that I have rounded some numbers. Comment if you face any issues